Grade 11Physical Chemistry

a sample of MgSO4 contain 8*10^20 oxygen atom find sulphur atom mass

MAYANK Agrawal
8 Years agoGrade 11

2 Answers

Adarsh
8 Years ago
Dear Mayank,
No of sulphur atom = ¼. No of oxygen atom= 2 x 1020 atom
Mass of sulphur in sample = 2 x 1020 x 32 / 6.023 x 10 23 = 10.62 x 10-3 gram
Thanks and Regards
Adarsh
Neerav Singla
6 Years ago
For every 4 atoms of oxygen, there is 1 atom of sulphur
So, for 1 atoms of oxygen, there will be ¼ atoms of sulphur
Acc. to the above relation for 8*1020 atoms there, will be 2*1020 atoms of sulphur
Mass of 2*1020 atoms of sulphur = 32*1.66*10-24*2*1020 g= 1.062*10-2g
Here, 1.66*10-24g is the unified mass unit.
 
Thank You